xx RotatingSymmetry180

RotatingSymmetry180

Erik Schnetter <schnetter@cct.lsu.edu>

2004-04-08

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11checkxboundaries10Scope: privatex01  BOOLEAN

11Description: Check the boundaries for poison

11               10  x01  Default: no

11

11poisonxboundaries 10Scope: privatex01  BOOLEAN

11Description: Fill the symmetry boundary with a poison value before the symmetry is applied

11               10  x01  Default: no

11

11poisonxvalue 10Scope: privatex01  INT

11Description: Integer value (0..255) used to poison new timelevels (with memset)

11 Range 10 x01 Default: 254
110:255  10Must fit into a byte. Use 0 for zero, 255 for nan, and e.g. 113 for a large value.

11

11verbose  10Scope: privatex01  BOOLEAN

11Description: Produce screen output while applying boundary conditions

11               10  x01  Default: no

11

7 Interfaces

General Implements: rotatingsymmetry90

Uses header: Slab.h

8 Schedule This section lists all the variables which are assigned storage by thorn CactusNumerical/RotatingSymmetry90. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxPARAMCHECK x rot90xchecktensortypes x check tensor type definitions for consistency x

11x11Language:11c
11 x 11Type: 11function

11

SymmetryRegister x rot90xregistersymmetry x register symmetry boundaries x

11x11Language:11c
11 x 11Type: 11function

11

BoundaryConditions x rot90xcomputelevelextent x compute extent of level x

11x11Before: 11rot90xapplybc
11 x 11Language:11c
11 x 11Options: 11level
11 x 11Type: 11function

11

BoundaryConditions x rot90xapplybc x apply 90 degree rotational symmetry boundary condition x

11x11Language:11c
11 x 11Type: 11function